OpenStudy (anonymous):

Find the missing length. Round to the nearest tenth, if necessary. (1 point) A)441 B)22.5 C)64 D)19.4

OpenStudy (anonymous):



OpenStudy (anonymous):

Pythagoras : 21^2 - 8^2 =?

OpenStudy (anonymous):

the answer it do in 2 min plzz

OpenStudy (anonymous):

\[a^2+b^2=c^2\]\[8^2+b^2=21^2\]

OpenStudy (anonymous):

x^2= 21^2 -8^2 x^2= 144-64 x^2=377 x= the square root of 377 x= 19.4

OpenStudy (anonymous):

come on, put a LITTLE work into it :/

OpenStudy (anonymous):

thank you faneaustin

OpenStudy (anonymous):

you're welcome